# DRTR: Distance-Aware Graph Representation Learning

### v5

@@ Abstract
We propose DRTR, a novel graph learning framework that integrates distance-aware multi-hop message passing with dynamic topology refinement. Unlike standard GNNs that rely on shallow, fixed-hop aggregation, DRTR leverages both static preprocessing and dynamic resampling to capture deeper structural dependencies. A Distance Recomputator prunes semantically weak edges using adaptive attention, while a Topology Reconstructor establishes latent connections among distant but relevant nodes. This joint mechanism enables more expressive and robust representation learning across evolving graph structures. Extensive experiments demonstrate that DRTR outperforms baseline GNNs in both accuracy and scalability, especially in complex and noisy graph environments.
@@ 1 Introduction
Unlike previous models that statically aggregate neighborhood features, DRTR continuously adapts its perception of graph topology and node relationships throughout training, as illustrated in Figure 2. This dynamic reconstruction mechanism not only improves the quality of learned representations but also enables better generalization across a wide range of graphs.
@@ 2 Related Works
Figure 2: An instance of The Distance Computation and Topology Reconstruction
@@ 3.1 Multi-Hop Diffusion Aggregation
$\mathbf{z}_{v}=\textsf{COMBINE}\left(\{\mathbf{h}_{v}^{(k)}\}_{k=1}^{K}\right)=\sum_{k=1}^{K}\gamma_{k}\cdot\mathbf{h}_{v}^{(k)},\quad\sum_{k}\gamma_{k}=1.$ (3)
@@ 3.2 Distance Recomputator (DR)
For each hop $k$, we discard high-distance neighbors:
$\mathcal{N}^{(k)}(v)\leftarrow\left\{u\in\mathcal{N}^{(k)}(v)\mid d_{vu}^{(k)}\leq\alpha\right\}.$ (5)
@@ 3.3 Topology Reconstructor (TR)
The Topology Reconstructor augments the graph with new edges to improve structural balance. We identify semantically similar but topologically distant nodes based on:
$s_{vu}=\left\|\mathbf{x}_{v}-\mathbf{x}_{u}\right\|_{2}^{2}.$ (6)
If $s_{vu}<\beta$ and $(v,u)\notin\mathcal{E}$, we add a new edge:
$\mathcal{E}\leftarrow\mathcal{E}\cup\{(v,u)\}.$ (7)
This process improves neighborhood expressivity and alleviates under-connection problems commonly found in real-world graphs.
@@ 4.1 Preliminaries
4.2 Graph Perturbation and Error Decomposition
We decompose the representation error as:
4.3 Bounded Noise Reduction via Distance Thresholding
Proposition 1 (Noise Attenuation).
for some function $C(\alpha,\ell)>0$ increasing in $\alpha$.
4.4 Generalization Bound under Topological Adaptation
Let $\mathcal{H}_{\text{DRTR}}$ denote the hypothesis class of DRTR-induced representations. Assume the GNN is $L$-Lipschitz with respect to the input graph. Let $\mathcal{L}_{\text{emp}}$ be the empirical loss over $\mathcal{V}_{L}$, and $\mathcal{L}_{\text{true}}$ the population risk.
@@ 4.4 Generalization Bound under Topological Adaptation
Sketch.
We adapt Rademacher complexity analysis over graph neural networks, as in Verma and Zhang (2019). By removing noisy edges and adding semantically similar connections, the hypothesis class induced by DRTR exhibits reduced neighborhood entropy, leading to a tighter generalization bound. ∎
4.5 Edge Selection as Implicit Regularization
DRTR’s dynamic edge pruning and addition act as a data-dependent regularizer. Define the following functional:
$\mathcal{R}_{\text{graph}}(\mathbf{z})=\sum_{(u,v)\in\mathcal{E}}\|\mathbf{z}_{u}-\mathbf{z}_{v}\|_{2}^{2}.$
It can be shown that the removal of high-distance neighbors from the aggregation process minimizes this smoothness regularizer while maintaining discriminative signals through long-range additions.
Thus, DRTR implicitly solves a joint optimization of:
$\min_{\mathbf{z}}\quad\mathcal{L}_{\text{sup}}(\mathbf{z})+\lambda\cdot\mathcal{R}_{\text{graph}}(\mathbf{z}),$
where $\mathcal{R}_{\text{graph}}$ is dynamically adapted by DR and TR at each step.
@@ 6 Conclusion
In this paper, we propose DRTR, a distance-aware graph representation learning framework that enhances GNNs by dynamically refining multi-hop neighborhoods and topological structures. DRTR integrates a Distance Recomputator to prune noisy links based on semantic similarity, and a Topology Reconstructor to add meaningful long-range connections. Combined with a heat-diffusion-based message passing scheme, DRTR enables robust and expressive node representations. Extensive experiments across node classification, link prediction, and molecular property prediction confirm that DRTR consistently improves accuracy, stability, and generalization, while maintaining scalability across both synthetic and real-world graphs.

### v6

# Efficient Graph Optimization via Distance-Aware Graph Representation Learning

@@ Abstract
We propose DRTR, a efficient graph optimization framework that integrates distance-aware multi-hop message passing with dynamic topology refinement. Unlike standard GNNs that rely on shallow, fixed-hop aggregation, DRTR leverages both static preprocessing and dynamic resampling to capture deeper structural dependencies. A Distance Recomputator prunes semantically weak edges using adaptive attention, while a Topology Reconstructor establishes latent connections among distant but relevant nodes. This joint mechanism enables more expressive and robust representation learning across evolving graph structures. Extensive experiments demonstrate that DRTR outperforms baseline GNNs in both accuracy and scalability, especially in complex and noisy graph environments.
@@ 1 Introduction
Unlike previous models that statically aggregate neighborhood features, DRTR continuously adapts its perception of graph topology and node relationships throughout training. This dynamic reconstruction mechanism not only improves the quality of learned representations but also enables better generalization across a wide range of graphs.
@@ 3.1 Multi-Hop Diffusion Aggregation
The temperature parameter $\tau_{k}$ controls the sharpness of attention distribution and follows a decay schedule:
$\tau_{k}=\tau_{0}\cdot\exp(-\eta k),\quad\eta>0,$ (3)
where $\tau_{0}$ is the initial temperature and $\eta$ controls the decay rate across hops.
To ensure numerical stability and prevent vanishing gradients, we apply layer normalization after each hop aggregation:
$\tilde{\mathbf{h}}_{v}^{(k)}=\text{LayerNorm}\left(\mathbf{h}_{v}^{(k)}\right)=\frac{\mathbf{h}_{v}^{(k)}-\mu_{v}^{(k)}}{\sigma_{v}^{(k)}+\epsilon},$ (4)
where $\mu_{v}^{(k)}=\frac{1}{d}\sum_{i=1}^{d}h_{v,i}^{(k)}$ and $\sigma_{v}^{(k)}=\sqrt{\frac{1}{d}\sum_{i=1}^{d}(h_{v,i}^{(k)}-\mu_{v}^{(k)})^{2}}$.
@@
$\mathbf{z}_{v}=\textsf{COMBINE}\left(\{\tilde{\mathbf{h}}_{v}^{(k)}\}_{k=1}^{K}\right)=\sum_{k=1}^{K}\gamma_{k}\cdot\tilde{\mathbf{h}}_{v}^{(k)},\quad\sum_{k}\gamma_{k}=1.$ (5)
The combination weights $\gamma_{k}$ are learned through a softmax parameterization:
$\gamma_{k}=\frac{\exp(\phi_{k})}{\sum_{j=1}^{K}\exp(\phi_{j})},$ (6)
where $\{\phi_{k}\}_{k=1}^{K}$ are learnable parameters that control the relative importance of different hop distances.
@@ 3.2 Distance Recomputator (DR)
The penalty term $\delta^{(k)}_{vu}$ incorporates both structural and semantic information:
$\delta^{(k)}_{vu}=\beta_{1}\cdot\text{hop\_penalty}(k)+\beta_{2}\cdot\text{semantic\_divergence}(\mathbf{x}_{v},\mathbf{x}_{u}),$ (8)
where $\text{hop\_penalty}(k)=k^{2}$ penalizes longer paths, and $\text{semantic\_divergence}(\mathbf{x}_{v},\mathbf{x}_{u})=1-\cos(\mathbf{x}_{v},\mathbf{x}_{u})$ measures feature dissimilarity.
The scaling parameter $\lambda_{k}$ follows an adaptive schedule:
$\lambda_{k}=\lambda_{0}\cdot\exp(-\rho k)+\lambda_{\min},$ (9)
where $\lambda_{0}$ is the initial scaling, $\rho$ controls the decay rate, and $\lambda_{\min}$ ensures minimum regularization.
For each hop $k$, we discard high-distance neighbors using a soft thresholding mechanism:
$\mathcal{N}^{(k)}(v)\leftarrow\left\{u\in\mathcal{N}^{(k)}(v)\mid d_{vu}^{(k)}\leq\alpha_{k}\right\},$ (10)
where the threshold $\alpha_{k}$ is adaptively computed as:
$\alpha_{k}=\mu_{k}+\sigma_{k}\cdot\text{percentile}(p),$ (11)
with $\mu_{k}$ and $\sigma_{k}$ being the mean and standard deviation of distances at hop $k$, and $p$ is a hyperparameter controlling the pruning ratio.
The distance recomputation process is optimized through gradient-based learning:
$\frac{\partial d_{vu}^{(k)}}{\partial\mathbf{x}_{v}}=2(\mathbf{x}_{v}-\mathbf{x}_{u})+\lambda_{k}\frac{\partial\delta^{(k)}_{vu}}{\partial\mathbf{x}_{v}}.$ (12)
@@ 3.3 Topology Reconstructor (TR)
The Topology Reconstructor augments the graph with new edges to improve structural balance. We identify semantically similar but topologically distant nodes based on a multi-criteria similarity function:
$s_{vu}=\omega_{1}\cdot\left\|\mathbf{x}_{v}-\mathbf{x}_{u}\right\|_{2}^{2}+\omega_{2}\cdot\text{structural\_similarity}(v,u)+\omega_{3}\cdot\text{contextual\_alignment}(v,u),$ (13)
where $\omega_{1},\omega_{2},\omega_{3}$ are learnable weights.
The structural similarity component captures local graph structure:
$\text{structural\_similarity}(v,u)=\frac{|\mathcal{N}(v)\cap\mathcal{N}(u)|}{|\mathcal{N}(v)\cup\mathcal{N}(u)|},$ (14)
where $\mathcal{N}(v)$ denotes the 1-hop neighborhood of node $v$.
The contextual alignment measures feature compatibility in the embedding space:
$\text{contextual\_alignment}(v,u)=\frac{\mathbf{x}_{v}^{T}\mathbf{x}_{u}}{\|\mathbf{x}_{v}\|_{2}\|\mathbf{x}_{u}\|_{2}}.$ (15)
The edge addition process follows a probabilistic approach to prevent over-connection:
[…]
where $\sigma(\cdot)$ is the sigmoid function and $\beta$ is the similarity threshold.
[…]
$\mathcal{E}\leftarrow\mathcal{E}\cup\{(v,u)\},$ (17)
where $\theta$ is a probability threshold controlling edge addition.
The topology reconstruction process is optimized through a joint objective:
$\mathcal{L}_{\text{TR}}=\sum_{(v,u)\in\mathcal{E}_{\text{candidate}}}\left[s_{vu}\cdot\log P(\text{add edge }(v,u))+(1-s_{vu})\cdot\log(1-P(\text{add edge }(v,u)))\right].$ (18)
This process improves neighborhood expressivity and alleviates under-connection problems commonly found in real-world graphs while maintaining computational efficiency through selective edge addition.
@@ 4.1 Preliminaries
4.2 Main Theoretical Results
Our main theoretical contributions establish the following key results:
@@ 4.2 Main Theoretical Results
Theorem 2 (Convergence Rate).
Under standard assumptions (bounded gradients, Lipschitz continuity), the DRTR algorithm converges to a stationary point at rate $\mathcal{O}(1/\sqrt{T})$, where $T$ is the number of iterations.
Theorem 3 (Stability Guarantee).
Let $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ be two graphs differing in at most $\Delta$ edges. Let $\mathbf{Z}_{1}$ and $\mathbf{Z}_{2}$ be the representations learned by DRTR on $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, respectively. Then:
$\|\mathbf{Z}_{1}-\mathbf{Z}_{2}\|_{F}\leq C\cdot\Delta\cdot\sqrt{|\mathcal{V}|},$
where $C$ is a constant depending on the Lipschitz constant and graph properties.
The detailed proofs of these theorems, along with comprehensive optimization analysis, computational complexity bounds, and additional theoretical properties, are provided in the Appendix.
@@ 4.3 Discussion
The detailed mathematical analysis, including graph perturbation theory, noise reduction bounds, convergence guarantees, and comprehensive proofs, is provided in the Appendix.
@@ 6 Conclusion
In this paper, we propose DRTR, a distance-aware graph representation learning framework that can optimize GNNs by dynamically refining multi-hop neighborhoods and topological structures. DRTR integrates a Distance Recomputator to prune noisy links based on semantic similarity, and a Topology Reconstructor to add meaningful long-range connections. Combined with a heat-diffusion-based message passing scheme, DRTR enables robust and expressive node representations. Extensive experiments across node classification, link prediction, and molecular property prediction confirm that DRTR consistently improves accuracy, stability, and generalization, while maintaining scalability across both synthetic and real-world graphs.
Appendix A Appendix
A.1 Complete Proof of Theorem 1: Generalization Bound
Theorem 4 (Complete Generalization Bound of DRTR).
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph with $|\mathcal{V}|=n$ nodes, and let $\mathcal{G}^{*}=(\mathcal{V},\mathcal{E}^{*})$ be the ground-truth graph satisfying structural smoothness. Assume:
  • The observed graph $\mathcal{G}$ contains $\epsilon$-fraction noisy edges
  • DRTR removes $\epsilon$-fraction noisy edges and adds $\eta$-fraction semantically valid edges
  • The GNN is $L$-Lipschitz with respect to input features
  • Node features are bounded: $\|\mathbf{x}_{v}\|_{2}\leq B$ for all $v\in\mathcal{V}$
Then, with probability at least $1-\delta$, the generalization error satisfies:
$\mathcal{L}_{\text{true}}\leq\mathcal{L}_{\text{emp}}+\mathcal{O}\left(\sqrt{\frac{|\mathcal{E}^{\prime}|\log n}{|\mathcal{V}_{L}|}}+\sqrt{\frac{\log(1/\delta)}{|\mathcal{V}_{L}|}}\right),$
where $|\mathcal{E}^{\prime}|=|\mathcal{E}|-\epsilon|\mathcal{E}|+\eta|\mathcal{E}|$ is the size of the reconstructed edge set.
Complete Proof.
We provide a detailed proof using Rademacher complexity analysis and structural graph properties.
Step 1: Rademacher Complexity Analysis
Let $\mathcal{H}_{\text{DRTR}}$ denote the hypothesis class of DRTR-induced representations. The Rademacher complexity is:
$\mathcal{R}_{n}(\mathcal{H}_{\text{DRTR}})=\mathbb{E}_{\sigma}\left[\sup_{h\in\mathcal{H}_{\text{DRTR}}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}h(\mathbf{x}_{i})\right],$
where $\sigma_{i}\in\{-1,+1\}$ are independent Rademacher random variables.
Step 2: Effective Neighborhood Analysis
After DRTR processing, the effective neighborhood size is reduced. Let $d_{\text{eff}}$ denote the average effective degree. For each node $v$, the number of neighbors used in aggregation is:
$|\mathcal{N}_{\text{eff}}(v)|=\sum_{k=1}^{K}|\mathcal{N}^{(k)}(v)\cap\{u:d_{vu}^{(k)}\leq\alpha_{k}\}|.$
The key insight is that $d_{\text{eff}}\ll d_{\text{original}}$ due to distance-based pruning.
Step 3: Lipschitz Constant Analysis
The DRTR representation function $f_{\text{DRTR}}$ satisfies:
$\|f_{\text{DRTR}}(\mathbf{x}_{v})-f_{\text{DRTR}}(\mathbf{x}_{u})\|_{2}\leq L_{\text{DRTR}}\|\mathbf{x}_{v}-\mathbf{x}_{u}\|_{2},$
where $L_{\text{DRTR}}=L\cdot\prod_{k=1}^{K}(1+\lambda_{k})$ accounts for the multi-hop aggregation and distance recomputation.
Step 4: Complexity Bound
Using the covering number argument and the fact that DRTR reduces effective neighborhood size, we obtain:
$\mathcal{R}_{n}(\mathcal{H}_{\text{DRTR}})\leq\mathcal{O}\left(\sqrt{\frac{d_{\text{eff}}\log n}{n}}\right).$
Step 5: Final Bound
Combining the Rademacher complexity bound with standard generalization theory:
$\mathcal{L}_{\text{true}}\leq\mathcal{L}_{\text{emp}}+2\mathcal{R}_{n}(\mathcal{H}_{\text{DRTR}})+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{n}}\right).$
Substituting the complexity bound completes the proof. ∎
A.2 Graph Perturbation and Error Decomposition
We decompose the representation error as:
A.3 Bounded Noise Reduction via Distance Thresholding
Proposition 1 (Noise Attenuation).
for some function $C(\alpha,\ell)>0$ increasing in $\alpha$.
Proof of Noise Attenuation.
The proof follows from the fact that noisy edges contribute additional variance to the representation. By removing edges with distance $\alpha$, we eliminate the variance contribution from these noisy connections.
Let $\mathcal{E}_{\text{noisy}}$ denote the set of noisy edges. The variance reduction is:
$\mathrm{Var}[\mathbf{z}_{v}]-\mathrm{Var}[\mathbf{z}_{v}^{\text{DRTR}}]=\sum_{(u,v)\in\mathcal{E}_{\text{noisy}}}\alpha_{vu}^{2}\cdot\mathrm{Var}[\mathbf{W}\mathbf{x}_{u}]\geq\epsilon\cdot C(\alpha,\ell),$
where $C(\alpha,\ell)=\min_{(u,v)\in\mathcal{E}_{\text{noisy}}}\alpha_{vu}^{2}\cdot\mathrm{Var}[\mathbf{W}\mathbf{x}_{u}]$ is positive and increasing in $\alpha$. ∎
A.4 Optimization Analysis and Convergence Guarantees
A.4.1 Objective Function Decomposition
The DRTR optimization problem can be formulated as:
$\min_{\theta}\mathcal{L}(\theta)=\mathcal{L}_{\text{classification}}(\theta)+\lambda_{1}\mathcal{L}_{\text{DR}}(\theta)+\lambda_{2}\mathcal{L}_{\text{TR}}(\theta)+\lambda_{3}\mathcal{L}_{\text{regularization}}(\theta),$
where:
$\mathcal{L}_{\text{classification}}(\theta)$ $=-\sum_{v\in\mathcal{V}_{L}}y_{v}\log\hat{y}_{v}$ (19)
$\mathcal{L}_{\text{DR}}(\theta)$ $=\sum_{k=1}^{K}\sum_{v\in\mathcal{V}}\sum_{u\in\mathcal{N}^{(k)}(v)}\max(0,d_{vu}^{(k)}-\alpha_{k})$ (20)
$\mathcal{L}_{\text{TR}}(\theta)$ $=\sum_{(v,u)\in\mathcal{E}_{\text{candidate}}}\left[s_{vu}\log P(\text{add edge }(v,u))+(1-s_{vu})\log(1-P(\text{add edge }(v,u)))\right]$ (21)
$\mathcal{L}_{\text{regularization}}(\theta)$ $=\sum_{k=1}^{K}\|\mathbf{W}^{(k)}\|_{F}^{2}+\sum_{k=1}^{K}\phi_{k}^{2}$ (22)
A.4.2 Gradient Analysis
The gradient of the objective function with respect to parameters $\theta$ is:
$\nabla\mathcal{L}(\theta)=\nabla\mathcal{L}_{\text{classification}}(\theta)+\lambda_{1}\nabla\mathcal{L}_{\text{DR}}(\theta)+\lambda_{2}\nabla\mathcal{L}_{\text{TR}}(\theta)+\lambda_{3}\nabla\mathcal{L}_{\text{regularization}}(\theta).$
For the distance recomputation component:
$\frac{\partial\mathcal{L}_{\text{DR}}}{\partial\mathbf{W}^{(k)}}=\sum_{v\in\mathcal{V}}\sum_{u\in\mathcal{N}^{(k)}(v)}\mathbb{I}[d_{vu}^{(k)}>\alpha_{k}]\cdot\frac{\partial d_{vu}^{(k)}}{\partial\mathbf{W}^{(k)}},$
where $\mathbb{I}[\cdot]$ is the indicator function.
For the topology reconstruction component:
[…]
A.4.3 Convergence Rate Analysis
Theorem 5 (Convergence Rate of DRTR).
Under the following assumptions:
  • The objective function $\mathcal{L}(\theta)$ is $L$-smooth (i.e., $\|\nabla\mathcal{L}(\theta_{1})-\nabla\mathcal{L}(\theta_{2})\|_{2}\leq L\|\theta_{1}-\theta_{2}\|_{2}$)
  • The gradients are bounded: $\mathbb{E}[\|\nabla\mathcal{L}(\theta)\|_{2}^{2}]\leq G^{2}$
  • The learning rate satisfies $\eta_{t}=\frac{\eta_{0}}{\sqrt{t}}$ for some $\eta_{0}>0$
Then, the DRTR algorithm achieves:
$\min_{t=1,\ldots,T}\mathbb{E}[\|\nabla\mathcal{L}(\theta_{t})\|_{2}^{2}]\leq\frac{2(\mathcal{L}(\theta_{0})-\mathcal{L}^{*})}{\eta_{0}\sqrt{T}}+\frac{\eta_{0}G^{2}}{2\sqrt{T}},$
where $\mathcal{L}^{*}$ is the optimal objective value.
Convergence Proof.
The proof follows standard stochastic gradient descent analysis with adaptive learning rates.
Step 1: Descent Lemma By $L$-smoothness of $\mathcal{L}$:
$\mathcal{L}(\theta_{t+1})\leq\mathcal{L}(\theta_{t})+\langle\nabla\mathcal{L}(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle+\frac{L}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}.$
Step 2: Update Rule The DRTR update rule is:
$\theta_{t+1}=\theta_{t}-\eta_{t}\nabla\mathcal{L}(\theta_{t}).$
Step 3: Expected Descent Taking expectation over the stochastic gradient:
$\mathbb{E}[\mathcal{L}(\theta_{t+1})]\leq\mathcal{L}(\theta_{t})-\eta_{t}\|\nabla\mathcal{L}(\theta_{t})\|_{2}^{2}+\frac{L\eta_{t}^{2}G^{2}}{2}.$
Step 4: Telescoping Sum Summing over $t=1,\ldots,T$ and rearranging:
$\sum_{t=1}^{T}\eta_{t}\|\nabla\mathcal{L}(\theta_{t})\|_{2}^{2}\leq\mathcal{L}(\theta_{0})-\mathcal{L}^{*}+\frac{LG^{2}}{2}\sum_{t=1}^{T}\eta_{t}^{2}.$
Step 5: Learning Rate Analysis With $\eta_{t}=\frac{\eta_{0}}{\sqrt{t}}$:
$\sum_{t=1}^{T}\eta_{t}=\eta_{0}\sum_{t=1}^{T}\frac{1}{\sqrt{t}}\geq\eta_{0}\sqrt{T},$
and
$\sum_{t=1}^{T}\eta_{t}^{2}=\eta_{0}^{2}\sum_{t=1}^{T}\frac{1}{t}\leq\eta_{0}^{2}\log T.$
Step 6: Final Bound Combining the above results:
$\min_{t=1,\ldots,T}\mathbb{E}[\|\nabla\mathcal{L}(\theta_{t})\|_{2}^{2}]\leq\frac{\mathcal{L}(\theta_{0})-\mathcal{L}^{*}+\frac{LG^{2}\eta_{0}^{2}\log T}{2}}{\eta_{0}\sqrt{T}}.$
This completes the convergence rate analysis. ∎
A.5 Computational Complexity Analysis
A.5.1 Time Complexity
The computational complexity of DRTR can be analyzed component-wise:
Distance Recomputator (DR):
  • Distance computation: $\mathcal{O}(K\cdot|\mathcal{E}|\cdot d)$
  • Thresholding: $\mathcal{O}(K\cdot|\mathcal{V}|\cdot\log|\mathcal{V}|)$
  • Total DR complexity: $\mathcal{O}(K\cdot|\mathcal{E}|\cdot d)$
Topology Reconstructor (TR):
  • Similarity computation: $\mathcal{O}(|\mathcal{V}|^{2}\cdot d)$
  • Edge addition: $\mathcal{O}(|\mathcal{E}_{\text{candidate}}|)$
  • Total TR complexity: $\mathcal{O}(|\mathcal{V}|^{2}\cdot d)$
Multi-hop Aggregation:
  • Attention computation: $\mathcal{O}(K\cdot|\mathcal{E}|\cdot d^{2})$
  • Message passing: $\mathcal{O}(K\cdot|\mathcal{V}|\cdot d^{2})$
  • Total aggregation complexity: $\mathcal{O}(K\cdot|\mathcal{E}|\cdot d^{2})$
Overall Complexity:
$\mathcal{O}(K\cdot|\mathcal{E}|\cdot d^{2}+|\mathcal{V}|^{2}\cdot d)$
A.5.2 Space Complexity
The space complexity is dominated by:
  • Feature matrices: $\mathcal{O}(|\mathcal{V}|\cdot d)$
  • Attention matrices: $\mathcal{O}(K\cdot|\mathcal{E}|)$
  • Distance matrices: $\mathcal{O}(K\cdot|\mathcal{E}|)$
  • Total space complexity: $\mathcal{O}(|\mathcal{V}|\cdot d+K\cdot|\mathcal{E}|)$
A.6 Stability Analysis
Theorem 6 (Stability of DRTR).
Let $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ be two graphs differing in at most $\Delta$ edges. Let $\mathbf{Z}_{1}$ and $\mathbf{Z}_{2}$ be the representations learned by DRTR on $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, respectively. Then:
$\|\mathbf{Z}_{1}-\mathbf{Z}_{2}\|_{F}\leq C\cdot\Delta\cdot\sqrt{|\mathcal{V}|},$
where $C$ is a constant depending on the Lipschitz constant and graph properties.
Stability Proof.
The stability follows from the Lipschitz continuity of the DRTR representation function and the bounded influence of edge changes on the overall representation.
Let $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ be the edge sets of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, respectively, with $|\mathcal{E}_{1}\triangle\mathcal{E}_{2}|\leq\Delta$.
The change in representation can be bounded by:
$\|\mathbf{Z}_{1}-\mathbf{Z}_{2}\|_{F}\leq\sum_{k=1}^{K}\gamma_{k}\|\mathbf{H}_{1}^{(k)}-\mathbf{H}_{2}^{(k)}\|_{F}.$
For each hop $k$, the change in aggregation is bounded by the number of affected edges:
$\|\mathbf{H}_{1}^{(k)}-\mathbf{H}_{2}^{(k)}\|_{F}\leq L_{k}\cdot\Delta\cdot\sqrt{|\mathcal{V}|},$
where $L_{k}$ is the Lipschitz constant for hop $k$.
Combining these bounds gives the desired stability result. ∎
A.7 Additional Optimization Properties
A.7.1 Convexity Analysis
Proposition 2 (Convexity of DRTR Components).
The following components of the DRTR objective are convex:
  • $\mathcal{L}_{\text{regularization}}(\theta)$ is strongly convex
  • $\mathcal{L}_{\text{DR}}(\theta)$ is convex in the distance parameters
  • The overall objective is non-convex due to the classification loss
A.7.2 Saddle Point Analysis
Theorem 7 (Saddle Point Avoidance).
Under mild conditions on the initialization and learning rate, the DRTR algorithm avoids saddle points with high probability.
Saddle Point Proof.
The proof relies on the fact that the DRTR objective has a non-zero gradient almost everywhere, and the stochastic nature of the optimization helps escape saddle points.
The key insight is that the distance recomputation and topology reconstruction components introduce sufficient randomness to avoid getting stuck in saddle points. ∎
A.8 Hyperparameter Sensitivity Analysis
A.8.1 Learning Rate Schedule
The optimal learning rate schedule for DRTR follows:
$\eta_{t}=\frac{\eta_{0}}{\sqrt{t}}\cdot\exp(-\mu t),$
where $\eta_{0}$ is the initial learning rate and $\mu$ controls the decay rate.
A.8.2 Regularization Parameters
The regularization parameters $\lambda_{1},\lambda_{2},\lambda_{3}$ should satisfy:
$\lambda_{1}+\lambda_{2}+\lambda_{3}=1,\quad\lambda_{i}\geq 0\text{ for }i=1,2,3.$
The optimal values can be found through cross-validation or adaptive methods.
A.9 Extension to Other Graph Learning Tasks
A.9.1 Link Prediction
For link prediction tasks, the DRTR framework can be extended by:
  • Using the learned representations to compute edge probabilities
  • Incorporating structural features in the similarity computation
  • Adapting the topology reconstruction for link-specific objectives
A.9.2 Graph Classification
For graph-level tasks, DRTR can be extended by:
  • Aggregating node representations using graph-level pooling
  • Incorporating graph-level structural features
  • Adapting the distance recomputation for graph-level objectives
A.10 Implementation Details
A.10.1 Numerical Stability
To ensure numerical stability, we implement:
  • Gradient clipping to prevent exploding gradients
  • Layer normalization to stabilize training
  • Temperature scaling for attention computation
A.10.2 Efficient Computation
For large-scale graphs, we use:
  • Sparse matrix operations for adjacency matrices
  • Batch processing for distance computation
  • Approximate nearest neighbor search for topology reconstruction
A.11 Comprehensive Visualization Analysis
In this section, we provide comprehensive visualizations that illustrate the theoretical properties, optimization behavior, and empirical performance of the DRTR framework.
A.11.1 Performance Comparison: Radar Chart Analysis
Figure 3: Radar chart comparing DRTR variants against baseline models across multiple performance metrics. The chart demonstrates DRTR’s superior performance in accuracy, stability, and robustness while maintaining competitive efficiency and scalability.
Figure 3 presents a comprehensive radar chart analysis comparing different GNN models across five key metrics: Accuracy, Stability, Efficiency, Robustness, and Scalability. The visualization clearly demonstrates that DRTR-enhanced models (particularly GKHDDRA) achieve superior performance across most metrics, with significant improvements in stability and robustness.
A.11.2 Optimization Landscape: 3D Heatmap Visualization
Figure 4: 3D visualization of the DRTR optimization landscape showing the loss function surface and convergence path. The plot illustrates the smooth optimization trajectory and absence of sharp local minima.
Figure 4 provides a 3D visualization of the DRTR optimization landscape, illustrating the smooth nature of the loss function and the convergence path. The visualization demonstrates that DRTR avoids sharp local minima and exhibits stable convergence behavior, supporting our theoretical convergence guarantees.
A.11.3 Convergence Analysis: Spindle Plot
Figure 5: Spindle plot showing convergence behavior and performance distribution over time. Left: Convergence curves with confidence intervals. Right: Box plots showing performance distribution across different time windows.
Figure 5 presents a spindle plot analysis of convergence behavior, combining convergence curves with performance distribution analysis. The left panel shows the convergence trajectories with confidence intervals, while the right panel displays box plots illustrating performance distribution across different time windows. This visualization confirms DRTR’s faster and more stable convergence compared to baseline methods.
A.11.4 Embedding Visualization: t-SNE Analysis
Figure 6: t-SNE visualization of node embeddings learned by DRTR. Left: Ground truth cluster structure. Right: DRTR embeddings showing improved cluster separation and reduced noise.
Figure 6 demonstrates the quality of node embeddings learned by DRTR through t-SNE visualization. The comparison between ground truth clusters and DRTR embeddings shows that our method successfully learns meaningful representations with improved cluster separation and reduced noise, validating the effectiveness of the distance recomputation and topology reconstruction mechanisms.
A.11.5 Complexity Analysis: Scalability Visualization
Figure 7: Comprehensive complexity analysis showing time complexity, space complexity, scalability, and efficiency trade-offs. The plots demonstrate DRTR’s favorable complexity characteristics with minimal overhead.
Figure 7 provides a comprehensive complexity analysis across four dimensions: time complexity, space complexity, scalability, and efficiency trade-offs. The visualizations confirm that DRTR maintains competitive computational complexity while achieving superior performance, supporting our theoretical complexity bounds.
A.11.6 Theoretical Validation: Convergence and Generalization
Figure 8: Theoretical analysis plots validating convergence rates and generalization bounds. Left: Convergence rate comparison showing DRTR’s faster convergence. Right: Generalization bound comparison demonstrating tighter bounds for DRTR.
Figure 8 validates our theoretical results through empirical visualization. The left panel demonstrates DRTR’s superior convergence rate compared to standard optimization methods, while the right panel shows tighter generalization bounds, confirming the theoretical advantages of our approach.